# Neural Scaling Laws for Deep Regression

### v1

@@ II.3 Scaling laws with the model size
Figure 6 shows the MSE test loss data from Fig. 4(a)–(c), plotted as a function of FCN model size at a fixed training dataset size. The test loss for $\theta$ generally decreases with increasing model size, but exhibits nonmonotonic, irregular patterns that do not follow a power-law scaling [Fig. 6(a)]. In contrast, the test loss for $J$ shows a more consistent, monotonic decreasing trend with increasing model size, spanning several orders of magnitude [Fig. 6(b)]. Notably, this reduction becomes more pronounced with larger dataset sizes, resulting in a more uniform decline toward the largest model size considered. Conversely, for smaller dataset sizes, the test loss tends to plateau at higher model sizes and, in some cases, shows a slight upturn—for example, in the curves for $N_{D}=224,448$. A similar decreasing trend, with substantial reduction at larger dataset sizes, is also observed for parameter $D$ [Fig. 6(c)]. However, in this case, the decrease is more modest and accompanied by notable fluctuations at small and large dataset sizes. Only intermediate dataset sizes display a monotonic decreasing trend. The fluctuations at the two largest dataset size ($N_{D}=57344,114688$) are likely due to the test loss approaching the limits of numerical precision.
Figure 7: Power-law scaling exponents $\alpha_{M}$ as a function of training dataset size. Each marker displays the fitted parameter of $\alpha_{M}$ from the MSE test loss data in Fig. 6(b)–(c) for $J$ and $D$, respectively. The training dataset size $N_{D}$ ranges from 224 to 114688. The dashed lines represent logarithmic fits of the form $\alpha_{M}=a_{M}\ln(N_{D})+b_{M}$, where $a_{M}$ and $b_{M}$ are fitting parameters.
Figure 7 shows the power-law scaling exponents $\alpha_{M}$, obtained from fitting the test loss data for parameters $J$ and $D$ in Fig. 6(b)–(c) to the relation $\epsilon\sim N_{M}^{-\alpha_{M}}$. The $\alpha_{M}$ values for $J$ increase significantly from $0.22\pm 0.05$ to $0.99\pm 0.12$ as the dataset size grows. These exponents are well described by a logarithmic fit $\alpha_{M}=a_{M}\ln N_{D}+b_{M}$, with fitted parameters $a_{M}=0.151\pm 0.008$ and $b_{M}=-0.59\pm 0.06$. In contrast, the $\alpha_{M}$ values for $D$ show only a weak increasing trend with the dataset size within a small range from $0.18\pm 0.05$ to $0.43\pm 0.10$, without significant variation or clear monotonicity. The fitted parameters in this case are $a_{M}=0.038\pm 0.015$ and $b_{M}=-0.01\pm 0.12$, with large uncertainties, indicating no consistent increase. The data for the two largest dataset sizes ($N_{D}=57344,114688$) do not conform well to the power-law fit for both $J$ and $D$ and are omitted from Fig. 7.
@@ III Discussion
In this study, we demonstrate the presence of a neural scaling law in deep regression through the widespread power-law relationships observed across several orders of magnitude in dataset size, model size, and diverse network architectures—including FCN, ResNet-18, and ViT. The large scaling exponents, reaching up to 2.3 with respect to dataset and model sizes, suggest that the accuracy of deep regression can be significantly improved simply by increasing data resources, or even more substantially by employing large models. The detailed scaling laws presented here offer valuable insights for designing deep regression models with predictable performance based on dataset size and network architecture, thereby mitigating resource-intensive trial-and-error experimentation. These findings are especially beneficial for applications in the physical sciences [32, 33, 34], such as condensed matter physics [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 25], high energy and particle physics [47, 48, 49] and astrophysics [50, 51, 52], where data often come in the form of simulated or experimental images, but data resources are frequently limited.
The emergence of substantial scaling exponents in generalization error exceeding two is noteworthy, especially since previous reports typically observe exponents less than one. For example, Hestness et al. found exponents in the range 0.07–0.35 for image classifications and language modeling [7]; Kaplan et al. reported an exponent of 0.095 for language modeling [18]; Henigham et al. observed exponents between 0.037 and 0.24 for image classifications [9]; Bahri et al. reported exponents ranging from 0.26 to 1.22 for image classifications [17]; and other studies have shown similar trends [6, 8, 11, 10, 12, 13, 14, 15, 16, 19, 20]. The first three studies used larger datasets and model sizes than those in this work, while the last study employed datasets and models of similar scale. This may explain the smaller exponents observed in the initial studies, with some overlap in exponents compared to our results and the last work. Additionally, network architecture appears to be a key factor; we found that FCNs tend to exhibit larger scaling exponents than ResNet-18 and ViT, a trend also observed in other studies. A detailed comparison and a comprehensive overview of the currently identified scaling laws are provided in the Supplementary Information.
@@ IV.1 Dataset generation
To generate magnetic domain images for twisted van der Waals magnets, the following Heisenberg spin model for twisted bilayer CrI3 [22] was employed:
@@
Here, $\bm{S}_{i}^{l}$ represents the spin operator at $i$ site on the top layer for $l=t$ or the bottom layer for $l=b$. The parameter $J>0$ represents intralayer ferromagnetic (FM) exchange interactions, $D>0$ represents single-ion anisotropy energy, and $J_{ij}^{\perp}$ represents interlayer exchange interactions. All these magnetic parameters are measured in meV units. The system is constructed on a commensurate moiré superlattice comprising two honeycomb lattice layers rotated relative to each other by a twist angle $\theta$. In total, the spin model is characterized by the four parameters $J$, $D$, $J_{ij}^{\perp}$, and $\theta$. The spin model is particularly relevant to the material CrI3 for the parameter values of $J\approx 2$ and $D\approx 0.2$ [53].
To obtain diverse magnetic domain images, parameter sets ($\theta$, $J$, $D$) were randomly sampled within the following ranges: $\theta$ from 1.01° to 3.89°, $J$ from 1 to 10, and $D$ from 0.01 to 0.3. The values of $J_{ij}^{\perp}$ were fixed, as determined by previous ab initio calculations [22]. Parameter sets corresponding to the FM state were excluded, since the FM state does not exhibit magnetic domains or variation with respect to parameter changes; excluding these states helps prevent potential overfitting [54]. In total, 162,782 parameter sets were used in our analysis. The magnetic ground states for each set were obtained through atomistic spin simulations. From each ground state, two magnetic domain images of size $100\times 100$ pixels were generated, each depicting out-of-plane magnetization in either the top or bottom layer. Consequently, our dataset comprises 325,564 paired images that illustrate the morphological variations of magnetic domains in twisted bilayer ferromagnets associated with the parameter changes. Examples of these images are shown in Fig. 2(a), where the parameters used are $J=1,2,3,4,5\,\text{meV}$, arranged from the first to fifth columns, and $\theta=1,2,3^{\circ}$, arranged from the first to third rows. Further details about the data generation process can be found in the Supplementary Information.
@@ IV.2 Deep learning algorithms
A test set of 20000 paired images was randomly selected from the total dataset of 162,782 paired images. The remaining data was bootstrapped to sizes of $N_{D}=256,512,\ldots,131072$, and each subset was divided into training and validation datasets with an 7/8 and 1/8 split, respectively. Pixel values of all images were normalized and standardized to enhance network performance. Each network was trained using the Adam optimizer for 200 epochs with early stopping patience set to 100. An adjustable learning rate was employed, starting at 0.001 and decreasing by a factor of 0.98 after each epoch. This training scheme was optimized to achieve the best performance, especially for the largest dataset, as detailed in the Supplementary Information. For each network architecture and dataset size, experiments were repeated approximately 20 times for FCN and ResNet, and around 10 times for ViT. All deep learning training and evaluation were conducted using PyTorch [55] on the Google Colab platform.
Fully connected network—Most employed FCNs consist of $n_{l}=3$ hidden layers, with each hidden layer containing $n_{n}=4,8,\ldots,512$ neurons. The input images’ pixel values were flattened and concatenated into a single feature vector of length $n_{i}=20{,}000$. The GELU activation function was used between all layers. The network outputted a single value corresponding to the Hamiltonian parameter via regression. The number of trainable parameters (weights and biases) is given by: $n_{n}\left[n_{i}+(n_{l}-1)n_{n}+1+n_{l}\right]+1\approx n_{i}n_{n}$, since $n_{i}\gg n_{n},n_{l},1$.
Residual network—The convolutional input layer of ResNet-18 was customized to consist of two input channels, instead of the standard three, to accommodate a pair of grayscale images of size $100\times 100$ pixels. The other hyperparameters of the input layer were kept unchanged. Additionally, the output layer was modified to produce a single scalar value for regression, as opposed to the original 1000-class classification. Consequently, the number of trainable parameters was reduced to 11,173,889 from the original 11,689,512. Due to these modifications, pretrained weights were not used; instead, the network was initialized randomly and trained from scratch.

### v2

@@ II.3 Scaling laws with the model size
Figure 6 shows the MSE test loss data from Fig. 4(a)–(c), plotted as a function of FCN model size at a fixed training dataset size. The test loss for $\theta$ generally decreases with increasing model size, but exhibits nonmonotonic, irregular patterns that do not follow a power-law scaling [Fig. 6(a)]. In contrast, the test loss for $J$ shows a more consistent, monotonic decreasing trend with increasing model size, spanning several orders of magnitude [Fig. 6(b)]. Notably, this reduction becomes more pronounced with larger dataset sizes, resulting in a more uniform decline toward the largest model size considered. Conversely, for smaller dataset sizes, the test loss tends to plateau at higher model sizes and, in some cases, shows a slight upturn—for example, in the curves for $N_{D}=224,448$. A similar decreasing trend, with substantial reduction at larger dataset sizes, is also observed for parameter $D$ [Fig. 6(c)]. However, in this case, the decrease is more modest and accompanied by notable fluctuations at small and large dataset sizes. Only intermediate dataset sizes display a monotonic decreasing trend. The fluctuations at the two largest dataset size ($N_{D}=57344,114688$) are likely due to the test loss approaching the limits of numerical precision approximately given by $1.9\times 10^{-7}$. This constraint is imposed by the 32-bit floating-point arithmetic used for the computation of the regressed parameters and the corresponding MSE test loss.
Figure 7: Power-law scaling exponents $\alpha_{M}$ as a function of training dataset size. Each marker displays the fitted parameter of $\alpha_{M}$ from the MSE test loss data in Fig. 6(b)–(c) for $J$ and $D$, respectively. The training dataset size $N_{D}$ ranges from 224 to 114688. The dashed lines represent logarithmic fits of the form $\alpha_{M}=a_{M}\log_{10}(N_{D}/10^{4})+b_{M}$, where $a_{M}$ and $b_{M}$ are fitting parameters.
Figure 7 shows the power-law scaling exponents $\alpha_{M}$, obtained from fitting the test loss data for parameters $J$ and $D$ in Fig. 6(b)–(c) to the relation $\epsilon\sim N_{M}^{-\alpha_{M}}$. The $\alpha_{M}$ values for $J$ increase significantly from $0.22\pm 0.05$ to $0.99\pm 0.12$ as the dataset size grows. These exponents are well described by a logarithmic fit, $\alpha_{M}=a_{M}\log_{10}(N_{D}/10^{4})+b_{M}$, with fitted parameters $a_{M}=0.347\pm 0.018$ and $b_{M}=0.799\pm 0.017$. In contrast, the $\alpha_{M}$ values for $D$ show only a weak increasing trend with the dataset size within a small range from $0.18\pm 0.05$ to $0.43\pm 0.10$, without significant variation or clear monotonicity. The fitted parameters in this case are $a_{M}=0.087\pm 0.035$ and $b_{M}=0.341\pm 0.032$, with large uncertainties, indicating no consistent increase. The data for the two largest dataset sizes ($N_{D}=57344,114688$) do not conform well to the power-law fit for both $J$ and $D$ and are omitted from Fig. 7.
@@ III Discussion
In this study, we demonstrate the presence of a neural scaling law in deep regression through the widespread power-law relationships observed across several orders of magnitude in dataset size, model size, and diverse network architectures—including FCN, ResNet-18, and ViT. The large scaling exponents, reaching up to 2.3 with respect to dataset and model sizes, suggest that the accuracy of deep regression can be significantly improved simply by increasing data resources, or even more substantially by employing large models. The detailed scaling laws presented here offer valuable insights for designing deep regression models with predictable performance based on dataset size and network architecture, thereby mitigating resource-intensive trial-and-error experimentation. These findings are especially beneficial for applications in the physical sciences [Dunjko2018, Carleo2019, Mehta2019], such as condensed matter physics [ohtsuki2016deep2d, Carrasquilla2017, PhysRevE.95.062122, ohtsuki2020drawing, Bedolla_2021, Lee2023, doi:10.1126/sciadv.abb0872, kwon2021magnetic, Lee2021, doi:10.1021/acsami.2c12848, Cadez2023MLFractalPhases, Bayo2025MLCondensedMatter, Lee_2024], high energy and particle physics [Radovic2018, Albertsson2018, Guest2018] and astrophysics [Hezaveh2017, Ntampaka2019, Reiman2020], where data often come in the form of simulated or experimental images, but data resources are frequently limited.
The emergence of substantial scaling exponents in generalization error exceeding two is noteworthy, especially since previous reports typically observe exponents less than one. For example, Hestness et al. found exponents in the range 0.07–0.35 for image classifications and language modeling [hestness2017deeplearningscalingpredictable]; Kaplan et al. reported an exponent of 0.095 for language modeling [kaplan2020scalinglawsneurallanguage]; Henigham et al. observed exponents between 0.037 and 0.24 for image classifications [henighan2020scaling]; Bahri et al. reported exponents ranging from 0.26 to 1.22 for image classifications [bahri2024explaining]; and other studies have shown similar trends [BAILLY2022106504, rosenfeld2020constructive, meir2020powerlaw, mühlenstädt2024dataneed2predicting, Figueroa2012, cho2016dataneededtrainmedical, 8237359, JMLR:v23:20-1111, NEURIPS2022_8c22e5e9, hoffmann2022trainingcomputeoptimallargelanguage, hutter2021learningcurvetheory, Defilippis2023DimensionFree, Paquette2024phases]. The first three studies used larger datasets and model sizes than those in this work, while the last study employed datasets and models of similar scale. This may explain the smaller exponents observed in the initial studies, with some overlap in exponents compared to our results and the last work. However, the absence of large scaling exponents in smaller dataset size regimes in the first three studies [hestness2017deeplearningscalingpredictable, kaplan2020scalinglawsneurallanguage, henighan2020scaling] challenges this explanation, suggesting that other factors (assigned regression tasks and specific scaling analysis methodologies) may also play significant roles. Additionally, network architecture appears to be a key factor; we found that FCNs tend to exhibit larger scaling exponents than ResNet-18 and ViT, a trend also observed in other studies. A detailed comparison and a comprehensive overview of the currently identified scaling laws are provided in the Supplementary Information.
@@ IV.1 Dataset generation
To generate magnetic domain images for twisted van der Waals magnets, the following Heisenberg spin model for twisted bilayer CrI3 [Kim2023] was employed:
@@
Here, $\bm{S}_{i}^{l}$ represents the spin operator at $i$ site on the top layer for $l=t$ or the bottom layer for $l=b$. The parameter $J>0$ represents intralayer ferromagnetic (FM) exchange interactions, $D>0$ represents single-ion anisotropy energy, and $J_{ij}^{\perp}$ represents interlayer exchange interactions. All these magnetic parameters are measured in meV units. The system is constructed on a commensurate moiré superlattice comprising two honeycomb lattice layers rotated relative to each other by a twist angle $\theta$. The twist angle $\theta$ determines the periodicity of $J_{ij}^{\perp}$, which is approximately given by $L\simeq a/\theta$, where $a$ is the lattice constant of the honeycomb lattice. In total, the spin model is characterized by the four parameters $J$, $D$, $J_{ij}^{\perp}$, and $\theta$. The spin model is particularly relevant to the material CrI3 for the parameter values of $J\approx 2$ and $D\approx 0.2$ [PhysRevX.8.041028].
To obtain diverse magnetic domain images, parameter sets ($\theta$, $J$, $D$) were randomly sampled within the following ranges: $\theta$ from 1.01° to 3.89°, $J$ from 1 to 10, and $D$ from 0.01 to 0.3. The values of $J_{ij}^{\perp}$ were fixed, as determined by previous ab initio calculations [Kim2023]. Parameter sets corresponding to the FM state were excluded, since the FM state does not exhibit magnetic domains or variation with respect to parameter changes; excluding these states helps prevent potential overfitting [Hawkins2004]. In total, 162,782 parameter sets were used in our analysis. The magnetic ground states for each set were obtained through atomistic spin simulations. From each ground state, two magnetic domain images of size $100\times 100$ pixels were generated, each depicting out-of-plane magnetization in either the top or bottom layer. Consequently, our dataset comprises 325,564 paired images that illustrate the morphological variations of magnetic domains in twisted bilayer ferromagnets associated with the parameter changes. Examples of these images are shown in Fig. 2(a), where the parameters used are $J=1,2,3,4,5\,\text{meV}$, arranged from the first to fifth columns, and $\theta=1,2,3^{\circ}$, arranged from the first to third rows. Further details about the data generation process can be found in the Supplementary Information.
@@ IV.2 Deep learning algorithms
A test set of 20000 paired images was randomly selected from the total dataset of 162,782 paired images. The remaining data was bootstrapped to sizes of $N_{D}=256,512,\ldots,131072$, and each subset was divided into training and validation datasets with an 7/8 and 1/8 split, respectively. Pixel values of all images were normalized and standardized to enhance network performance. Each network was trained using the Adam optimizer for 200 epochs with early stopping patience set to 100. An adjustable learning rate was employed, starting at 0.001 and decreasing by a factor of 0.98 after each epoch. This training scheme was optimized to achieve the best performance, especially for the largest dataset, as detailed in the Supplementary Information. For each network architecture and dataset size, experiments were repeated approximately 20 times for FCN and ResNet, and around 10 times for ViT. All deep learning training and evaluation were conducted using PyTorch [PyTorch2019] on the Google Colab platform.
Fully connected network—Most employed FCNs consist of $n_{l}=3$ hidden layers, with each hidden layer containing $n_{n}=4,8,\ldots,512$ neurons. The input images’ pixel values were flattened and concatenated into a single feature vector of length $n_{i}=20{,}000$. The GELU activation function was used between all layers. The network outputted a single value corresponding to the Hamiltonian parameter via regression. The number of trainable parameters (weights and biases) – the model size $N_{M}$, is given by: $N_{M}=n_{n}\left[n_{i}+(n_{l}-1)n_{n}+1+n_{l}\right]+1\approx n_{i}n_{n}$, since $n_{i}\gg n_{n},n_{l},1$.
Residual network—The convolutional input layer of ResNet-18 was customized to consist of two input channels, instead of the standard three, to accommodate two grayscale images of size $100\times 100$ pixels, each for the top and bottom layer domain structures, respectively. The other hyperparameters of the input layer were kept unchanged. Additionally, the output layer was modified to produce a single scalar value for regression, as opposed to the original 1000-class classification. Consequently, the number of trainable parameters was reduced to 11,173,889 from the original 11,689,512. Due to these modifications, pretrained weights were not used; instead, the network was initialized randomly and trained from scratch.